The sun has a mass of 1.99 × 10^33 kg If Neptune is 4.50 × 10^12m from the sun. What is the acceleration of Neptune due to the sun’s gravity? Give your answer in Scientific Notation.


1
Expert's answer
2021-02-12T16:17:44-0500

We can find the acceleration of Neptune due to the Sun’s gravity from the formula:


"a=\dfrac{GM_{Sun}}{R^2},""a=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot1.99\cdot10^{33}\ kg}{(4.5\cdot10^{12}\ m)^2}=6.55\cdot10^{-3}\ \dfrac{m}{s^2}."